OpenStudy (anonymous):

A and B are events such that P(A+B)=3/4, P(AB)=1/4 and P(A')=2/3. What is P(B)?

OpenStudy (anonymous):

Firstly find P(A).. P(A) = 1 - P(A') can you find P(A)..??

OpenStudy (anonymous):

yup...then??

OpenStudy (anonymous):

Tell me what is it..

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Use: \[\large \color{green}{P(A + B) = P(A) + P(B) - P(AB)}\]

OpenStudy (anonymous):

p(a)=1-2/3=1/3 p(b)=p(a+b)+p(ab)-p(a)=3/4 +1/4 - 1/3=2/3

OpenStudy (anonymous):

@matricked why are you giving solution in one go???

OpenStudy (anonymous):

Don't look at the solution do it yourself @Mitul

OpenStudy (anonymous):

thank u....i was confused if P(A U B) and P(A+B) were same...@waterineyes thank u very much

OpenStudy (anonymous):

See: P(A or B) = P(A+B) P(A and B) = P(A*B) = P(AB) \(\color{green}{AND}\) represents Multiplication.. \(\color{green}{OR}\) represents Addition..

OpenStudy (anonymous):

yup :)